Year Price of Good A Quantity of Good A Price of Good B Quantity of Good B Price of Good C Quantity of Good C

Base Year

2013

$35 100 $95 200 $75 150
2014 $40 125 $97 275 $78 175
2015 $42 150 $98 280 $80 200
2016 $45 250 $100 290 $85 225

Using the table above and 2013 as the base year, compute Nominal GDP, Real GDP, the GDP Deflator for each year (2013 through 2016). Compute the inflation rate for each of the following years (2014, 2015 and 2016).
Answer #1

Nominal GDP of 2013 = 35 * 100 + 95 * 200 + 75*150 = 33750

Real GDP of 2013 = 35 * 100 + 95 * 200 + 75*150 = 33750

GDP Deflator = 33750 / 33750 * 100 = 100

Nominal GDP of 2014 = 40 * 125 + 97 * 275 + 78*175 = 45325

Real GDP of 201 = 35 * 125 + 95 * 275 + 75*175 = 43625

GDP Deflator of 2014 = 45325 / 43625 * 100 = 103.90

Inflation in 2014 = (103.90 - 100) / 100 * 100 = 3.90%

Nominal GDP of 2015 = 42 * 150 + 98 * 280 + 80 * 200 = 49740

Real GDP of 2015 = 35 * 150 + 95 * 280 + 75*200 = 46850

GDP Deflator for 2015 = 49740 / 46850 * 100 = 106.17

Inflation in 2015 = (106.17 - 103.90) / 103.90 * 100 = 2.18%

Nominal GDP of 2016 = 45 * 250 + 100 * 290 + 85 * 225 = 59375

Real GDP of 2015 = 35 * 250 + 95 * 290 + 75*225 = 53175

GDP Deflator for 2015 = 59375 / 53175 * 100 = 111.66

Inflation in 2015 = (111.66 - 106.17) / 106.17 * 100 = 5.17%